# Surrogate Modeling of 3D Rayleigh-Bénard Convection with Equivariant Autoencoders

### v1

@@ Abstract
The use of machine learning for modeling, understanding, and controlling large-scale physics systems is quickly gaining in popularity, with examples ranging from electromagnetism over nuclear fusion reactors and magneto-hydrodynamics to fluid mechanics and climate modeling. These systems—governed by partial differential equations—present unique challenges regarding the large number of degrees of freedom and the complex dynamics over many scales both in space and time, and additional measures to improve accuracy and sample efficiency are highly desirable. We present an end-to-end equivariant surrogate model consisting of an equivariant convolutional autoencoder and an equivariant convolutional LSTM using $G$-steerable kernels. As a case study, we consider the three-dimensional Rayleigh-Bénard convection, which describes the buoyancy-driven fluid flow between a heated bottom and a cooled top plate. While the system is E(2)-equivariant in the horizontal plane, the boundary conditions break the translational equivariance in the vertical direction. Our architecture leverages vertically stacked layers of $D_{4}$-steerable kernels, with additional partial kernel sharing in the vertical direction for further efficiency improvement. Our results demonstrate significant gains both in sample and parameter efficiency, as well as a better scaling to more complex dynamics, that is, larger Rayleigh numbers. The accompanying code is available under https://github.com/FynnFromme/equivariant-rb-forecasting.
@@ 1 Introduction
The ability to perform fast and efficient simulations of large scale physics systems governed by partial differential equations (PDEs) is of vital importance in many areas of science and engineering. As in almost any other area, machine learning plays an increasingly important role, where scenarios of great interest are real-time prediction, uncertainty quantification, optimization and control. Application areas include weather forecasting (Kurth et al., 2023) and climate modeling (Vlachas et al., 2018), aerodynamics and fluid mechanics (Brunton et al., 2020), combustion (Ihme et al., 2022), or the plasma in nuclear fusion reactors (Kates-Harbeck et al., 2019). PDE-governed systems often exhibit complex or chaotic behavior over a vast range of scales in both space and time. Along with the very large number of degrees of freedom (after discretization using, e.g., finite elements), this renders the resulting time series particularly challenging for surrogate modeling, especially in multi-query contexts such as prediction and control (Bieker et al., 2020). Fortunately, many dynamical systems evolve on a low-dimensional manifold (e.g., an attractor), which allows for dimensionality reduction techniques and surrogate modeling. As linear approximation techniques such as proper orthogonal decomposition (POD) (Sirovich, 1987) tend to break down once the dynamics become more complex—as is the case for turbulent flows—nonlinear variants such as autoencoders have recently become more and more important in the physics modeling community (Nikolopoulos et al., 2022; Francés-Belda et al., 2024). However, the resulting reduced spaces are less interpretable and the training is much more data-hungry and sensitive to hyperparameter tuning.
The goal of this paper is to include known symmetries in the surrogate modeling process of complicated 3D physics simulations via autoencoders. Studying a prototypical convection or climate system described as Rayleigh-Bénard convection, our contributions are the following (cf. Fig. 1 for a sketch):
@@
  • We develop an end-to-end equivariant architecture for the prediction of 3D time-dependent PDEs consisting of an autoencoder and an LSTM for time series prediction in latent space. To preserve the symmetry, we omit latent space flattening but preserve the original 3D structure of the problem.
@@
  • The system is equivariant under rotations, reflections and translations (i.e., the symmetry group $E(n)$), but only in the horizontal plane due to the buoyancy. We introduce an efficient $G$-steerable autoencoder architecture that respects the symmetry group $\mathbb{Z}^{2}\rtimes D_{4}$ in the horizontal plane (the subgroup of $E(2)$ with shifts on a grid and discrete 90-degree rotations).
  • Equivariance in the vertical direction is only broken by the boundary conditions (the equations themselves are equivariant under vertical translations). We incorporate additional local kernel sharing in the vertical direction and study its effects on performance and efficiency.
@@
  • We demonstrate high accuracy at a compression rate $>98\%$, while saving one order of magnitude in both trainable parameters and training data compared to non-symmetric architectures.
Figure 1: Architecture overview. (Left-right) An initial snapshot $s_{t}$ is encoded via our E(2) equivariant 3D autoencoder to the latent representation $z_{t}$; $z_{t}$ is evolved forward in time to $\tilde{z}_{t+1}$ via our equivariant LSTM; $\tilde{z}_{t+1}$ is decoded via our E(2) equivariant 3D decoder to yield a predicted next snapshot $\tilde{s}_{t+1}$.
@@ 2 Related work
In recent years, a large number of works have appeared that address the task of surrogate modeling of PDE systems. Most of these approaches rely on the identification of a low-dimensional latent space for instance via POD (Soucasse et al., 2019), autoencoders (Pandey et al., 2022; Akbari et al., 2022; de Sousa Almeida et al., 2023) or simply random measurements (Gao et al., 2025) (even though the latter does not explicitly address Rayleigh-Bénard convection). Alternatively, the direct prediction of the full state can be accelerated using linear methods such as the Koopman operator framework (Klus et al., 2020; Markmann et al., 2024) or physics-informed machine learning (Karniadakis et al., 2021; Clark Di Leoni et al., 2023; Hammoud et al., 2023).
The exploitation of symmetries has recently become increasingly popular, and it is now often referred to under the umbrella term geometric deep learning (Bronstein et al., 2021). Most of the literature in this area is until now related to classical learning tasks such as image classification (Cohen and Welling, 2016; Esteves et al., 2018a, b; Weiler and Cesa, 2019; Bronstein et al., 2021; Weiler et al., 2025). However, equivariant architectures have been developed for various other tasks such as the analysis of graph-structured data (e.g., molecules (Wu et al., 2021)), or for the prediction of PDEs (Jenner and Weiler, 2022; Zhdanov et al., 2024; Harder et al., 2024b). Equivariant autoencoder architectures for dimensionality reduction were proposed in, e.g., Kuzminykh et al. (2018); Guo et al. (2019); Huang et al. (2022), see also Hao et al. (2023); Yasuda and Onishi (2023) for applications to fluid flows. Further examples of equivariant learning of PDEs were presented in various contexts such as Koopman operator theory and Dynamic Mode Decomposition (Salova et al., 2019; Baddoo et al., 2023; Harder et al., 2024a; Peitz et al., 2025), as well as reinforcement learning (Vignon et al., 2023; Vasanth et al., 2024; Peitz et al., 2024; Jeon et al., 2024).
@@ 3.1 Symmetries
If a group action is defined on $X$, one can obtain an action on the space of functions of the form $\phi:X\rightarrow\mathbb{R}^{m}$ by introducing $(\rho(g)\phi)(x):=\phi(g^{-1}\cdot x)$. Here, we have already denoted the action as a representation, see (Cohen and Welling, 2017), as it is linear.
@@
[…]
This is an example of a translation equivariant operation, as $[\rho(\tau)f_{\mathsf{in}}]*\psi=\rho(\tau)[f_{\mathsf{in}}*\psi]$ for $\tau\in\mathbb{R}^{n}$ (Cohen and Welling, 2016).
Steerable convolutions. In steerable CNNs, we define feature spaces in such a manner that the respective feature fields $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{C}$ are steerable (Cohen and Welling, 2017; Weiler and Cesa, 2019). This means that for each $n$-dimensional input $x\in\mathbb{R}^{n}$, each $C$-dimensional feature $f(x)\in\mathbb{R}^{C}$ transforms under the group action of a given group $G$ (e.g., $O(n)$)—the translational equivariance is automatically satisfied when the parameters of the learnable kernel $\psi$ are position-independent (Weiler et al., 2021). In particular, this ensures that the transformation of vector-valued features transforms their orientation according to the group action. As a consequence, all CNN architectures that are constructed using $G$-steerable kernels are equivariant with respect to the group $H=(\mathbb{R}^{n},+)\rtimes G$. In Weiler and Cesa (2019), it was shown that for this property to hold, a kernel $\psi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{C_{\mathsf{out}}\times C_{\mathsf{in}}}$ has to satisfy the steerability constraint
@@
For a network to be equivariant, the kernel constraint (1) has to hold for all combinations of $\rho_{\mathsf{in}}$ and $\rho_{\mathsf{out}}$. Instead, it is shown in Weiler and Cesa (2019) that a much simpler approach is to introduce a change of basis $Q$, by which any $\rho$ can be decomposed into the direct sum of its irreducible representations (irreps), i.e., $\rho=Q^{-1}\left[\oplus_{i\in I}\psi_{i}\right]Q$. One can thus replace (1) by individual constraints on the irreps. Since we have $G\leq O(2)$, $G$ is norm preserving, such that the kernel constraint can further be reformulated in terms of a Fourier series expansion of the kernel, ultimately resulting in a set of constraints on the Fourier coefficients, cf. Weiler and Cesa (2019); Weiler et al. (2021) for detailed derivations and discussions. We will make heavy use of this approach in our architecture, both in terms of the autoencoder and the LSTM for time series prediction.
@@ 3.2 Rayleigh-Bénard convection
Rayleigh-Bénard convection describes the flow between two flat plates. The bottom plate is heated, while the top plate is cooled, which induces buoyancy forces that cause the fluid to move in convection rolls, cf. Figs. 1 or 2 for illustrations. Depending on the driving force—the temperature difference, encoded by the dimensionless Rayleigh number $Ra$—the system is deterministic at first, then becomes increasingly complex and turbulent for larger $Ra$ (Pandey et al., 2018).

### v2

@@ Abstract
The use of machine learning for modeling, understanding, and controlling large-scale physics systems is quickly gaining in popularity, with examples ranging from electromagnetism over nuclear fusion reactors and magneto-hydrodynamics to fluid mechanics and climate modeling. These systems—governed by partial differential equations—present unique challenges regarding the large number of degrees of freedom and the complex dynamics over many scales both in space and time, and additional measures to improve accuracy and sample efficiency are highly desirable. We present an end-to-end equivariant surrogate model consisting of an equivariant convolutional autoencoder and an equivariant convolutional LSTM using $G$-steerable kernels. As a case study, we consider the three-dimensional Rayleigh-Bénard convection, which describes the buoyancy-driven fluid flow between a heated bottom and a cooled top plate. While the system is E(2)-equivariant in the horizontal plane, the boundary conditions break the translational equivariance in the vertical direction. Our architecture leverages vertically stacked layers of $D_{4}$-steerable kernels, with additional partial kernel sharing in the vertical direction for further efficiency improvement. We demonstrate significant gains in sample and parameter efficiency, as well as a better scaling to more complex dynamics.
@@ 1 Introduction
The ability to perform fast and efficient simulations of large scale physics systems governed by partial differential equations (PDEs) is of vital importance in many areas of science and engineering. As in almost any other area, machine learning plays an increasingly important role, where scenarios of great interest are real-time prediction, uncertainty quantification, optimization and control. Application areas include weather forecasting (Kurth et al., 2023) and climate modeling (Vlachas et al., 2018), aerodynamics and fluid mechanics (Brunton et al., 2020), combustion (Ihme et al., 2022), or the plasma in nuclear fusion reactors (Kates-Harbeck et al., 2019). PDE-governed systems often exhibit complex or chaotic behavior over a vast range of scales in both space and time. Along with the very large number of degrees of freedom (after discretization using, e.g., finite elements), this renders the resulting time series particularly challenging for surrogate modeling, especially in multi-query contexts such as prediction and control (Bieker et al., 2020). Fortunately, many dynamical systems evolve on a low-dimensional manifold (e.g., an attractor), which allows for dimensionality reduction techniques and surrogate modeling. As linear approximation techniques such as proper orthogonal decomposition (POD) (Sirovich, 1987) tend to break down once the dynamics become more complex—as is the case for turbulent flows—nonlinear variants such as autoencoders have recently become more and more important in the physics modeling community (Nikolopoulos et al., 2022; Francés-Belda et al., 2024). However, the resulting reduced spaces are less interpretable and the training is much more data-hungry and sensitive to hyperparameter tuning. The goal of this paper is to include known symmetries in the surrogate modeling process of complicated 3D physics simulations via autoencoders. Studying a prototypical convection or climate system described as Rayleigh-Bénard convection, our contributions are the following (cf. Fig. 1 for a sketch):
@@
  • Whereas most papers are concerned with 2D in space, we develop an end-to-end equivariant architecture for the much more challenging prediction of 3D time-dependent PDEs, consisting of an autoencoder and an LSTM for time series prediction in latent space. To preserve the symmetry, we omit latent space flattening but preserve the original 3D structure of the problem.
@@
  • The system is equivariant under rotations, reflections, and translations (i.e., the symmetry group $E(n)$), but only in the horizontal plane due to the buoyancy. We introduce an efficient $G$-steerable autoencoder architecture that respects the symmetry group $\mathbb{Z}^{2}\rtimes D_{4}$ in the horizontal plane (the subgroup of $E(2)$ with shifts on a grid, reflections, and discrete 90-degree rotations).
  • Additional local kernel sharing in the vertical direction further improves the parameter efficiency.
  • We demonstrate that for large-scale systems, a separate training of encoding and time stepping can be computationally much more efficient than end-to-end training.
@@
  • We demonstrate high accuracy at a compression rate $98\%$, while saving one order of magnitude in both trainable parameters and training data compared to non-symmetric architectures.
@@ 2 Related work
Figure 1: Architecture overview. An initial snapshot $s_{t}$ is encoded via our E(2) equivariant 3D autoencoder to the latent representation $z_{t}$; $z_{t}$ is evolved forward in time to $\tilde{z}_{t+1}$ via our equivariant LSTM; $\tilde{z}_{t+1}$ is decoded via our E(2) equivariant 3D decoder to yield a predicted next snapshot $\tilde{s}_{t+1}$.
In recent years, a large number of works have appeared on surrogate modeling of PDE systems. Many approaches rely on the identification of a low-dimensional latent space for instance via POD (Soucasse et al., 2019) orautoencoders (Pandey et al., 2022; Akbari et al., 2022; de Sousa Almeida et al., 2023). Alternatively, the direct prediction of the full state can be accelerated using linear methods such as the Koopman operator framework (Klus et al., 2020; Markmann et al., 2024; Azencot et al., 2020; Nayak et al., 2025) or physics-informed machine learning (Karniadakis et al., 2021; Clark Di Leoni et al., 2023; Hammoud et al., 2023). In that latter category, neural operators (Li et al., 2021; Lu et al., 2021; Kovachki et al., 2023; Goswami et al., 2023; Straat et al., 2025) are very prominent, and have been demonstrated to show great performance on a large number of systems. Finally, there have been great advances in the deep learning area as well, most prominently using U-Nets (Gupta & Brandstetter, 2023; Lei & Li, 2025), transformer architectures (Gao et al., 2024; Holzschuh et al., 2025b) and generative frameworks such as GANs (Chen et al., 2020) or diffusion models (Holzschuh et al., 2025a; Bastek et al., 2025; Li et al., 2025; Oommen et al., 2025).
The exploitation of symmetries has recently become increasingly popular, and it is now often referred to under the umbrella term geometric deep learning (Bronstein et al., 2021). Most of the literature in this area is until now related to classical learning tasks such as image classification (Cohen & Welling, 2016; Esteves et al., 2018a; b; Weiler & Cesa, 2019; Bronstein et al., 2021; Weiler et al., 2025). However, equivariant architectures have been developed for various other tasks such as the analysis of graph-structured data (e.g., molecules (Wu et al., 2021)), or for the prediction of PDEs (Jenner & Weiler, 2022; Zhdanov et al., 2024; Harder et al., 2024b). Equivariant autoencoder architectures for dimensionality reduction were proposed in, e.g., Kuzminykh et al. (2018); Guo et al. (2019); Huang et al. (2022), see also Hao et al. (2023); Yasuda & Onishi (2023) for applications to fluid flows. Further examples of equivariant learning of PDEs were presented in various contexts such as Koopman operator theory and Dynamic Mode Decomposition (Salova et al., 2019; Baddoo et al., 2023; Harder et al., 2024a; Peitz et al., 2025), as well as reinforcement learning (Vignon et al., 2023; Vasanth et al., 2024; Peitz et al., 2024; Jeon et al., 2024).
@@ 3.1 Symmetries
If a group action is defined on $X$, one can obtain an action on the space of functions of the form $\phi:X\rightarrow\mathbb{R}^{m}$ by introducing $(\rho(g)\phi)(x):=\phi(g^{-1}\cdot x)$. Here, we have already denoted the action as a representation, see (Cohen & Welling, 2017), as it is linear.
@@
$f_{\mathsf{out}}(x)=[f_{\mathsf{in}}*\psi](x)=\int_{\mathbb{R}^{n}}\psi(x-y)f_{\mathsf{in}}(y)\,dy$
This is an example of a translation equivariant operation, as $\rho(\tau)f_{\mathsf{in}}]*\psi=\rho(\tau)[f_{\mathsf{in}}*\psi]$ for $\tau\in\mathbb{R}^{n}$ (Cohen & Welling, 2016).
@@
Steerable convolutions. In steerable CNNs, we define feature spaces in such a manner that the respective feature fields $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{C}$ are steerable (Cohen & Welling, 2017; Weiler & Cesa, 2019). This means that for each $n$-dimensional input $x\in\mathbb{R}^{n}$, each $C$-dimensional feature $f(x)\in\mathbb{R}^{C}$ transforms under the group action of a given group $G$ (e.g., O(n)O(n))—the translational equivariance is automatically satisfied when the parameters of the learnable kernel $\psi$ are position-independent (Weiler et al., 2021). In particular, this ensures that the transformation of vector-valued features transforms their orientation according to the group action. As a consequence, all CNN architectures that are constructed using $G$-steerable kernels are equivariant with respect to the group $H=(\mathbb{R}^{n},+)\rtimes G$. In Weiler & Cesa (2019), it was shown that for this property to hold, a kernel $\psi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{C_{\mathsf{out}}\times C_{\mathsf{in}}}$ has to satisfy the steerability constraint
@@
For a network to be equivariant, the kernel constraint (1) has to hold for all combinations of $\rho_{\mathsf{in}}$ and $\rho_{\mathsf{out}}$. Instead, it is shown in Weiler & Cesa (2019) that a much simpler approach is to introduce a change of basis $Q$, by which any $\rho$ can be decomposed into the direct sum of its irreducible representations (irreps), i.e., $\rho=Q^{-1}\left[\oplus_{i\in I}\psi_{i}\right]Q$. One can thus replace (1) by individual constraints on the irreps. Since we have $\leq O(2), GG$ is norm preserving, such that the kernel constraint can further be reformulated in terms of a Fourier series expansion of the kernel, ultimately resulting in a set of constraints on the Fourier coefficients, cf. Weiler & Cesa (2019); Weiler et al. (2021) for detailed derivations and discussions. We will make heavy use of this approach in our architecture, both in terms of the autoencoder and the LSTM for time series prediction.
@@ 3.2 Rayleigh-Bénard convection
Rayleigh-Bénard convection describes the flow between two flat plates. The bottom plate is heated, while the top plate is cooled, which induces buoyancy forces that cause the fluid to move in convection rolls, cf. Figs. 1 or 2 for illustrations. Depending on the driving force—the temperature difference, encoded by the dimensionless Rayleigh number RaRa—the system is deterministic at first, then becomes increasingly complex and turbulent for larger $Ra$ (Pandey et al., 2018). More details about the specific PDE and the numerical simulations can be found in Appendix A.1. A more detailed discussion on the system’s symmetries—including a proof—can be found in Appendix A.2. Moreover, a list of other systems with broken symmetries can be found in Appendix A.3.
4 Methods
4.1 3D steerable convolution on E(2)
As discussed in Appendix A.2, the 3D Rayleigh-Bénard convection is $E(2)$-equivariant in the horizontal plane, which we enforce by steerable convolutions, while introducing height-dependent kernels that adapt to the varying dynamics at different heights.
4.1.1 Steerable convolution
The system state $s$ consists of both scalar temperature and vector-valued velocity fields. Under transformation of these fields by $\tau g\in E(2)$, the temperature field $T:\mathbb{R}^{3}\rightarrow\mathbb{R}$ transforms as $T(x)\mapsto 1\cdot T(g^{-1}(x-\tau))$, whereas the velocity-field $u:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ transforms as $u(x)\mapsto g\cdot u(g^{-1}(x-\tau))$. Note that the velocity vectors are themselves transformed via $g$ to preserve their orientation as the field is transformed (Cohen & Welling, 2017; Weiler & Cesa, 2019).
Thus, a three-dimensional $O(2)$-steerable kernel can be constructed as a stack of 2D $O(2)$-steerable kernels. For these, Weiler & Cesa (2019) have solved the steerability constraint as well as for important subgroups such as $C_{4}$ and $D_{4}$. This result has been applied in our implementation to efficiently design height-dependent kernels utilizing the PyTorch-based library escnn. [^1] https://github.com/QUVA-Lab/escnn For computational reasons, we restrict our implementation to the subgroup $H=(\mathbb{Z}^{2},+)\rtimes D_{4}<E(2)$, as this optimally corresponds to our data on a rectangular grid. Thus, we will from now on consider discretized kernels and feature fields on $\mathbb{Z}^{3}$. The case of continuous rotations according to $O(2)$ would require interpolation between grid points (see, e.g., Esteves et al. (2018a)). In our experiments, this resulted in inferior performance compared to the grid-consistent 90-degree-rotations and flips.
Within our framework, we use various types of feature fields. Both the input to the AE-encoder and the output of the decoder are composed of the scalar field $T$ and the vector field $u$. All intermediate representations, however, make use of regular feature fields, which transform under the regular representation by permuting the channels. Steerable convolutions between regular feature fields are equivalent to regular group convolutions (Cohen & Welling, 2016), which apply the same kernel in every orientation $g\in G$, resulting in the $|G|$-dimensional regular feature fields.
4.1.2 3D convolutions with height-dependent kernels
Height-dependent features—such as temperature or velocity patterns—play a critical role in modeling the system. As a result, applying the same kernel across all heights—as would be the case for a regular 3D CNN—is insufficient to capture the system’s vertical dynamics. To address this limitation, we modify the conventional 3D convolutions by learning height-dependent steerable kernels. While this ensures horizontal parameter sharing, we allow the convolution operation to adapt to the distinct features at each height, ensuring that the vertical structure is captured effectively. For computing the output feature map’s value at position $x=(x_{1},x_{2},x_{3})\in\mathbb{Z}^{3}$, the input $f:\mathbb{Z}^{3}\rightarrow\mathbb{R}^{C_{\mathsf{in}}}$ is convolved with the height-dependent kernel $\psi_{x_{3}}:\mathbb{Z}^{3}\rightarrow\mathbb{R}^{C_{\mathsf{out}}\times C_{\mathsf{in}}}$ via $f*\psi](x)=\sum_{y\in\mathbb{Z}^{3}}\psi_{x_{3}}(x-y)f(y)$.
Local vertical parameter sharing. Although the Rayleigh-Bénard system does not exhibit global vertical translation equivariance, it approximately maintains this property within a local neighborhood (see Appendix A.2). This suggests that features at a given height $x_{3}$ are locally correlated with features at vertical positions within a $2k+1$-sized neighborhood $\mathcal{N}({x}_{3})=\{{x}_{3}-k,\ldots,{x}_{3}+k\}$. This approximate local equivariance in the vertical direction can be exploited by choosing a smaller number of channels in each layer, but then applying these learned kernels across the local neighborhood $\mathcal{N}$, thereby again increasing the number of output channels, cf. Fig. 3 for a sketch. This parameter reduction leads to a more efficient and scalable model, rendering it suitable for simulating larger-scale systems. As described in Appendix B.2, 3D convolutions with height-dependent kernels and local vertical parameter sharing can be efficiently implemented by wrapping 2D (steerable) convolutions.
4.2 Equivariant autoencoder
The decoder applies a series of convolutions, followed by activation functions. Pooling operations are replaced by upsampling. To preserve continuity while upsampling, we employ trilinear interpolation. Throughout the entire framework (CAE + LSTM), padding is applied to the convolutional layers to ensure that the spatial dimensions of the data are only altered by pooling and upsampling operations. For the horizontal dimensions, circular padding is used to match the periodic boundary conditions of the Rayleigh-Bénard system, while vertical dimensions use zero padding. Note that without padding, the decoder would require transposed convolutions (Dumoulin & Visin, 2018) to reverse the dimensional changes induced by convolutions without padding.
4.3 Equivariant LSTM
Long short-term memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) extend conventional recurrent neural networks (RNNs) for processing sequential data by an additional cell state $c_{t}$. While the hidden state $h_{t}$ stores currently relevant information for predicting the next time step, the cell state $c_{t}$ captures long-term information over long sequences. The content of $c_{t}$ is regulated by the input gate $i_{t}$, which controls which information is added, and the forget gate $f_{t}$, which controls how much information is discarded. The output gate $o_{t}$ determines which information is passed from $h_{t}$ to $c_{t}$. To preserve the spatial structure of the latent space $z$, we use convolutional LSTMs (Shi et al., 2015) that replace the fully connected layers of standard LSTMs with, in our case, equivariant convolutions and also introduce equivariant convolutions in peephole connections, see B.1 for details as well as a proof that the architecture is equivariant.
5 Experiments
We evaluate the performance of our end-to-end equivariant framework against a baseline model using standard, non-steerable 3D convolutions, with a particular focus on the autoencoder. In addition, we assess the long-term forecasting capabilities of our approach in comparison to 3D Fourier Neural Operators (FNOs) (Li et al., 2021) and 3D U-Nets (Ronneberger et al., 2015; Çiçek et al., 2016) with the same number of parameters, cf. Appendix B.3 for details.
Datasets. We generated a dataset of 100 randomly initialized 3D Rayleigh-Bénard convection simulations with $Ra=2500$ and $Pr=0.7$, standardized to zero mean and unit standard deviation. The dataset was split into 60 training, 20 validation, and 20 test simulations, each with 400 snapshots in the time interval $t\in[100,300]$ at a step size of 0.5 and a spatial resolution of $N=48\times 48\times 32$. For long-term forecasting evaluation, we also created an additional dataset of 20 simulations, each containing 1800 snapshots over $t\in[100,1000]$.
Architecture. The CAE encoder and decoder each consist of six convolutional layers with a kernel size of three for the last encoder and first decoder layer, and five for the other layers, applying ELU nonlinearities pointwise. [^2] All hidden layers in our model transform under the regular representation, which preserves equivariance since it commutes with pointwise nonlinearities (Weiler & Cesa, 2019). Pooling and upsampling are applied after the encoder and decoder’s first, third, and fifth layers, respectively. The latent space has dimensions $M=6\times 6\times 4$ and $C_{\mathsf{latent}}=32$ channels, resulting in a compression to 1.56% of the original size. Channels roughly double after each pooling operation in the encoder, with the decoder reversing this. Over our experiments, we vary the channel sizes. For a fair comparison, our main models (both standard and equivariant) have approximately the same number of 3.6M parameters. The decoder-only LSTM with one layer of convolutional LSTM cells and an additional convolutional layer for output prediction uses a kernel size of 3. The number of channels is chosen such that the LSTM has approximately 3.7M parameters.
Training. CAE training is performed in the standard self-supervised manner on a set of snapshots, where the desired decoder output is the original input. The LSTM was then trained on the latent representations to predict the 50 latent states following the provided sequence of 25 states, with the loss being computed on the latent space. This two-step training approach effectively separates encoding and forecasting, avoiding significant performance drops and speeding up forecasting training by a factor of 20. See Appendix C.1 for training details, and Appendix C.2 for an in-depth discussion of the two-step training approach and the limitations of end-to-end training.
5.1 Results
Figure 4: CAE performances. Results are shown with varying numbers of parameters, amounts of training samples, compression ratios, and Rayleigh numbers RaRa. Each plot shows a variation over one of the parameters while keeping the others fixed. The main model has 3.6 million parameters ($3^{\mathsf{rd}}$ point), 24,000 training samples ($4^{\mathsf{th}}$ point), 1.56% latent size ($3^{\mathsf{rd}}$ point) and $Ra=2500$ ($1^{\mathsf{st}}$ point).
Autoencoder. Fig. 4 gives an overview of our experiments to compare the equivariant and non-equivariant autoencoders with respect to parameter efficiency, data efficiency, compression capabilities, and scalability to more complex dynamics, i.e., larger $Ra$. Equivariance leads to a significant improvement in reconstruction accuracy, with the RMSE decreasing by 42%, from 0.069 $\pm$ 0.00072 to 0.04 $\pm$ 0.00093, where the mean and standard deviation were computed across three models with randomly initialized weights. A comparison to $C_{4}$-equivariant convolutions shows an RMSE of 0.046 $\pm$ 0.00164, showing that both rotations and reflections are relevant for the improved performance.
  1. (i)
    Parameter efficiency: The equivariant model with only 900,000 parameters outperformed the non-equivariant model with 10.8 million parameters, meaning that we obtain an improvement by more than one order of magnitude.
  2. (ii)
    Sample efficiency: The $D_{4}$ model achieved the same performance as the non-equivariant model when trained on just one-eighth of the training samples, representing a significant reduction in the amount of data required for effective learning, making it particularly valuable in data-limited scenarios or when simulations/experiments are expensive.
  3. (iii)
  4. (iv)
Long-term forecasting. We next focus on the evaluation of the long-term forecasting capabilities of our end-to-end equivariant architecture. The model is provided with an input sequence of 50 snapshots and then predicts the subsequent 500 future states in an autoregressive manner, i.e., using its own predictions as input for the next time step.
Figure 5: Forecasting RMSE averaged across three runs. FNO and U-Net baselines were restricted to 5 steps due to computational cost, while our model scales efficiently to 50 steps. For reference, we also include the $D_{4}$-steerable model trained with 5 steps.
Fig. 5 shows the median RMSE over time for both our equivariant and non-equivariant models, which was averaged over three separately trained models. The equivariant model consistently outperforms the non-equivariant model by a near-constant margin of approximately 0.05 RMSE across the entire forecast horizon. This indicates that most of the performance gains stem from improved latent representation learned by the equivariant autoencoder, which provides a more stable and informative basis for prediction.
Our method consistently outperforms both FNO and U-Net baselines across all horizons. With equal training horizons (5 steps), it already matches or exceeds their short-term accuracy. The decisive advantage, however, lies in the ability to efficiently train with much longer horizons: while extending FNOs and U-Nets beyond 5 autoregressive steps during training becomes prohibitively expensive (see Table 1 in Appendix C.1), our latent-space approach scales to 50 steps without difficulty, since forecasting is performed directly in latent space. This leads to substantially improved long-term forecasts. For completeness, we also report the performance of our model trained with only 5 steps, which still surpasses both baselines at short horizons, although the FNO slightly outperforms our approach at longer horizons.
These results highlight a central advantage of our two-step training strategy: it enables efficient training with long autoregressive horizons in latent space, achieving both superior long-term performance and significantly lower computational cost compared to those baselines.
Finally, we note that the advantages of our surrogate extend beyond comparisons with neural baselines. Appendix C.3 provides a detailed discussion of its benefits over classical PDE solvers, in particular inference speed, scalability to multi-query settings, and end-to-end differentiability.
6 Conclusion
Our equivariant CAE plus LSTM architecture for efficient surrogate modeling of 3D Rayleigh-Bénard convection consists of horizontally $E(2)$-equivariant kernels that are vertically stacked. Additional local sharing in the vertical direction allows us to increase the number of channels without requiring additional parameters, which further adds to the computational efficiency. We have demonstrated significant advantages in terms of the accuracy and both data and parameter efficiency.
6.1 Limitations and outlook
Some points we have not yet discussed or addressed, but believe to be promising for future research:
  • In principle, end-to-end training could be superior, as the latent representation is tailored to the dynamics. Since this proved to be challenging for LSTMs, we believe that a consecutive finetuning phase is more promising. Instead, one could consider linear latent dynamics based on the Koopman operator (Harder et al., 2024a; Azencot et al., 2020).
  • In the control setting, equivariant models are very helpful to improve the efficiency, for instance of world models in the context of reinforcement learning; surrogate modeling for RL of PDEs was studied in, e.g., Werner & Peitz (2024), but a combination with equivariant RL as proposed in van der Pol et al. (2020) has not been investigated.
  • Incorporating kernels that explicitly depend on control parameters (e.g., Rayleigh number) may improve adaptability and generalization across different flow regimes, including those with qualitative changes such as bifurcations.
Code
Acknowledgments
Appendix A Additional details on the Rayleigh–Bénard system and its symmetries
A.1 Boussinesq approximation and numerical solution